OpenStudy (anonymous):

                My son has $18 in his bank. How much more money does he need to save in order to buy the video game that he wants that costs $45?                                 18 + x = 45, with x representing the amount of money he needs to save.

OpenStudy (lgbasallote):

and the question is...?

OpenStudy (ganpat):

18 + x = 45, find for x.. subtract 18 from both sides.. x = 45 - 18 x = 27 $ more need to be saved..